find three consecutive odd integers such that the sum of the middle and the largest inteer is 21 more than the smallest integer
Answer by macston(5194)

.
S=smallest integer; M=middle integer=S+2; L=largest integer=S+4
.
M+L=S+21
S+2+S+4=S+21
2S+6=S+21
S=15
M=S+2=15+2=17
L=S+4=15+4=19
ANSWER: The integers are 15, 17, and 19.
.
CHECK:
M+L=S+21
17+19=15+21
36=36